OpenStudy (anonymous):

2.As you are leaving the community, you notice another new community just down the street. Because you are in the area, you decide to inquire about it. a.The sales representative here tells you they also have two floor plans available, but they only have 38 homes still for sale. Write an equation that illustrates the situation. Use x and y to denote floor plan one and floor plan two respectively. b.The representative tells you that floor plan one sells for $175,000 and floor plan two sells for $200,000. She also mentions that all the available houses combined are worth $7,200,000. Write an

OpenStudy (anonymous):

. Write an equation that illustrates this situation. Use the same variables you used in Part a. c. Use elimination to determine how many houses are available in each floor plan. Explain how you arrived at your answer

OpenStudy (anonymous):

175000x + 200000y = 7200000 x+y=38 y=38-x 175000x+200000(38-x)=7200000 175000x+7600000-200000x=7200000 7600000-25000x=7200000 -25000x=-400000 x=16 38-16=y y=22

OpenStudy (anonymous):

22 floor plan 2 18 floor plan 1

OpenStudy (anonymous):

Do you need explanation Rachel?

OpenStudy (anonymous):

I am looking them over so I can see which answer is for which question...and yes I may need some explanation

OpenStudy (anonymous):

ok, use group chat in right bottom corner or here if you want. standing by.

OpenStudy (anonymous):

ok thank you just give me a sec.

OpenStudy (anonymous):

a. x+y=38 b. 175000x + 200000y = 7200000 then c. would be the rest... 175000x+200000(38-x)=7200000 175000x+7600000-200000x=7200000 7600000-25000x=7200000 -25000x=-400000 x=16 38-16=y y=22

OpenStudy (anonymous):

all i did was solve for x in x+y=38 and place it in other equation

OpenStudy (anonymous):

solved for y^